The U. S. Environmental Protection Agency (EPA) sets limits on healthful levels of air pollutants. The maximum level that the EPA considers safe for lead air pollution is 1.5 μg/m3. If your lungs were filled with air containing this level of lead, how many lead atoms would be in your lungs? (Assume a total lung volume of 5.40 L .)
Answer #1

Conversion Factor

1m^3 = 1000 L

Mass of Lead in our Lungs = 1.5 ug/m^3 * 5.40L * 1m^3/1000L = 8.1 * 10^(-9) gms

Molar mass of Lead = 207.2

Number of moles of Lead = mass of Lead/molar mass

=> 8.1 * 10^(-9)/207.2

=> 3.90 * 10^(-11) moles

1 mole contains 6.023 * 10^(23) atoms

Hence number of lead atoms

=> 3.90 * 10^(-11) * 6.023 * 10^(23)

=> 2.349 * 10^(13) atoms